# Computing High-dimensional Confidence Sets for Arbitrary Distributions

### v1

@@ Abstract
Our main result is an algorithm that finds a confidence set whose volume is $\exp(\tilde{O}(d^{2/3}))$ factor competitive with the optimal ball having the desired coverage. It is surprisingly simple and also extends to finding confidence sets competitive against unions of $k$ balls, and other variants. The algorithm is improper (it outputs an ellipsoid). Combined with our computational intractability result for proper learning balls within an $\exp(\tilde{O}(d^{1-o(1)}))$ approximation factor in volume, our results provide an interesting separation between proper and (improper) learning of confidence sets.
@@ 1 Introduction
where $\operatorname*{\mathrm{vol}}(S)$ denotes the volume or Lebesgue measure of the set $S$. The minimum of (1) is known as the generalized quantile function (Einmahl and Mason, 1992; Polonik, 1997), an important quantity in statistical inference that plays a central role in a number of problems including estimating density level sets (Garcia et al., 2003), distribution support estimation (Schölkopf et al., 2001) and conformal prediction (Gao et al., 2025). A set containing at least $\delta$, say $0.9$, probability mass generalizes the idea of confidence intervals from the one-dimensional setting to a high-dimensional setting.
Solving (1) leads to a number of interesting applications in statistical inference. On the one hand, finding a set that contains a large probability mass of the distribution can be useful as a high-dimensional analogue of support estimation (Schölkopf et al., 2001). On the other hand, as $\delta\to 0$, the maximizer of (1) can be used as a mode estimator  (Sager, 1978, 1979). Algorithms that learn confidence sets for arbitrary distributions can also be used as a tool for estimating the uncertainty in the predictions of black-box machine learning models, as in conformal prediction (Gammerman et al., 1998; Vovk et al., 2005). Finally, algorithms for this task have to be robust to outliers and contamination present in the the remaining $(1-\delta)$ fraction of the data, and gives a different notion of robustness that is non-parametric and may be of independent interest in robust estimation (Rousseeuw, 1984, 1985; Van Aelst and Rousseeuw, 2009). There is also a deep and rich body of work on robust statistics, where the remaining $1-\delta$ fraction can be arbitrary outliers, but these works often make stronger model assumptions about the $\delta$ fraction of points that are “inliers”  (Huber, 1964; Diakonikolas and Kane, 2023). See Section 1.5 for several applications in statistics like conformal prediction, goodness-of-fit, robust estimation and testing unimodality.
@@
Consider the setting where there exists a good confidence set belonging to a family $\mathcal{C}$ of finite VC-dimension. When independent samples from the distribution $\mathcal{D}$ are available, one can solve the empirical version of (1) with polynomial sample complexity and establish approximate volume optimality as long as the concept class $\mathcal{C}$ has finite VC-dimension (Polonik, 1999; Scott and Nowak, 2005). This is related to the (agnostic) PAC learning setting in classification where the confidence sets $S$ can be identified with classifiers belonging to the concept class $\mathcal{C}$. This leads to the following broad class of learning problems:
@@
If there exists a small confidence set $C^{*}\in\mathcal{C}$ with (slightly more than) the required coverage $\delta$, then the algorithm will output a set $S$ with coverage $\delta$ and whose volume is competitive with $C^{*}$. The set $S$ is not required to belong to the family $\mathcal{C}$. This is analogous to the standard setting in PAC learning, where the learning algorithm may output a classifier that does not belong to the concept class $\mathcal{C}$ (see e.g., Shalev-Shwartz and Ben-David, 2014). If the algorithm outputs a confidence set $S$ that belongs to $\mathcal{C}$, we will call it a proper learner of confidence sets (analogous to proper PAC learning). When we want to be more explicit, the term improper learner may be used for an algorithm that does not necessarily output a set belonging to the concept class $\mathcal{C}$. While there is a simple algorithm that is sample efficient (but computationally inefficient) when the VC-dimension of $\mathcal{C}$ is bounded (by a polynomial), the algorithmic problem becomes challenging in high dimensions.
@@
Consider the basic setting where $\mathcal{C}$ is the set of $\ell_{2}$ balls (this has VC-dimension at most $d$). It is NP-hard to find the smallest set in $\mathcal{C}$ that covers a $\delta$ fraction of points from an arbitrary set (Shenmaier, 2015). [^1] On the other hand, finding the minimum volume ball that encloses all the given $n$ points is polynomial time solvable. The famous work of Badoiu, Har-Peled and Indyk on coresets for minimum volume enclosing balls (Badoiu et al., 2002; Agarwal et al., 2004) gives a proper learning algorithm that runs in polynomial time and is $\Gamma=1+O(1/\log d)$-competitive i.e., a $\exp\big{(}O(d/\log d)\big{)}$ factor approximation in volume. To the best of our knowledge, the work of Badoiu et al. (2002) remains the state-of-the-art guarantee for learning confidence sets competitive with Euclidean balls in $d$ dimensions. We focus on Euclidean balls and unions of Euclidean balls in the rest of the paper.
@@ 1.1 Our Results
Our main algorithmic result is an algorithm that is $\Gamma=1+O(d^{-1/3})$ competitive against Euclidean balls.
@@
There is a polynomial time algorithm that for any target coverage $\delta\in(0,1)$ and coverage slack $\gamma\in(0,1)$, when given $n=\Omega(d^{2}/\gamma^{2})$ samples drawn i.i.d. from an arbitrary distribution $\mathcal{D}$, finds with high probability a set $S\subset\mathbb{R}^{d}$ that is $\Gamma=\left(1+O_{\gamma,\delta}\big{(}d^{-1/3+o(1)}\big{)}\right)$ competitive, i.e.,
@@
$\operatorname*{\mathrm{vol}}(S)^{1/d}\leq\operatorname*{\mathrm{vol}}(B^{\star})^{1/d}\cdot\left(1+d^{-1/3+o(1)}\cdot O\left(\frac{1}{\delta}+\frac{1}{\sqrt{\gamma\delta}}\right)\right).$
@@
Please see Section 4 for more details. To compare against existing methods, consider the setting when $\delta,\gamma$ are constants. A straightforward solution that considers balls around the sample points that covers at least $\delta$ fraction of the data points achieves a factor of $\Gamma=2$ (due to the triangle inequality). [^2] In fact this strategy achieves a $\Gamma=2$ factor approximation (with $\textrm{vol}^{1/d}$) for any set $\mathcal{C}$ that consists of scalings and translations of a fixed convex shape. However, it is NP-hard to obtain a $(2-o(1))$ factor approximation even for the $\ell_{\infty}$ ball (Shenmaier, 2015). Algorithms based on the powerful technique of coresets for minimum volume enclosing balls due to  Badoiu et al. (2002) gives $\Gamma=1+O(1/\log d)$-competitive. In terms of the volume, the $\Gamma=1+d^{-1/3}$ factor corresponds to volume approximation of
$\Gamma^{d}=\big{(}1+O_{\delta,\gamma}(d^{-1/3+o(1)})\big{)}^{d}=\exp\left(O_{\gamma,\delta}\big{(}d^{2/3+o(1)}\big{)}\right),$
which corresponds to a substantial improvement from the $\exp\big{(}O_{\delta,\gamma}(d/\log d)\big{)}$ volume approximation factor in existing work (Badoiu et al., 2002).
@@
Let $\delta\in(0,1),\gamma\in(0,1),k\in\mathbb{N}$ be any constants. There is a polynomial time algorithm that for target coverage $\delta\in(0,1)$ and coverage slack $\gamma\in(0,1)$ when given $n=\Omega(kd^{2}/\gamma^{2})$ samples drawn i.i.d. from an arbitrary distribution $\mathcal{D}$, finds with high probability a set $S\subset\mathbb{R}^{d}$ that is $\Gamma=\left(1+O_{\gamma,\delta}\big{(}d^{-1/3+o(1)}\big{)}\right)$ competitive; more precisely, it satisfies $\mathbf{P}_{y\sim\mathcal{D}}\left[y\in S\right]\geq\delta$, and
$\operatorname*{\mathrm{vol}}(S)^{1/d}\leq\mathrm{vol}(C^{\star}_{k})^{1/d}\left(1+O_{k,\delta}\left(d^{-1/3+o(1)}\right)\right)\cdot\left(\frac{O(\log(k/\gamma))}{\gamma}\right)^{1/d}$
@@ 1.2 Proper Learning: Separation Result and Algorithms
For any constant $\varepsilon>0$, assuming $P\neq NP$ there is no polynomial time algorithm that given samples from an arbitrary distribution over $\mathbb{R}^{d}$ that outputs balls as confidence sets that are $\Gamma\leq(1+\frac{1}{d^{\varepsilon}})$ competitive with balls.
See Theorem 6.1 for a more formal statement. The above theorem follows from a simple reduction that amplifies the existing NP-hardness result due to (Shenmaier, 2015). In Theorem 6.3, we also show computational intractability (assuming the Small Set Expansion hypothesis of Raghavendra and Steurer (2010)) that shows a similar quantitative hardness, even when the coverage can be smaller by an arbitrary constant factor. Combined with our algorithmic result in Theorem 1.1, this shows a formal separation between proper and improper learning for a natural geometric learning problem.
@@
When we are required to output a confidence set from $\mathcal{C}$, the algorithm in Theorem 1.1 can also be used to output a ball with a worse competitive ratio (which is expected due to Theorem 1.3). When we make no assumptions about the distribution $\mathcal{D}$, our algorithm can output a ball that is $\Gamma=1+\widetilde{O}(1/\sqrt{\log d})$ competitive. See Theorem 3.1 in Section 3 for a formal statement of the theorem. This guarantee is slightly worse than the guarantee from the coreset-based algorithm which achieves $\Gamma=1+\widetilde{O}(1/\log d)$ (Badoiu et al., 2002). However, in the worst-case instances of the algorithm (and the hard instances in Theorem 1.3), the points sampled from $\mathcal{D}$ that lie inside the optimal ball $B^{\star}$ effectively lie on a lower-dimensional subspace.
@@
Given a set of $n$ points $Y\subseteq\mathbb{R}^{d}$, and $\delta,\gamma\in(0,1)$, such that there is an unknown ball $B^{\star}=B(c^{\star},R^{\star})$ with $Y^{\star}=B^{\star}\cap Y$ satisfying $|Y^{\star}|\geq\delta|Y|$, and $Y^{\star}$ is $\beta(R^{\star})^{2}/d$-isotropic for some $\beta\geq 0$. That is, if $\mu^{\star}$ is the mean of the points $Y^{\star}$, then
@@
$\operatorname*{\mathrm{vol}}(\widehat{B})^{1/d}\leq\operatorname*{\mathrm{vol}}(B^{\star})^{1/d}\left(1+O\left((\gamma^{-1/2}+\delta^{-1/2})\frac{\beta^{1/2}}{d^{1/2}}\right)\right),$
@@
In the above Theorem 1.4, $\Sigma_{Y^{\star}}$ is a $d\times d$ PSD matrix with $\mathrm{tr}(\Sigma_{Y^{\star}})\leq(R^{\star})^{2}$, since $Y^{\star}\subseteq B^{\star}$. The isotropicity condition requires the maximum eigenvalue $\lambda_{\max}(\Sigma_{Y^{\star}})\leq\beta(R^{\star})^{2}/d$ for some $\beta>0$. Thus every $Y^{\star}$ satisfies the isotropic condition with $\beta\leq d$. When the points of $Y^{\star}$ are very spread out, this corresponds to a setting when $\beta\leq 1$ (all the eigenvalues are equal). In this case, the theorem guarantees to find a ball with $\Gamma=1+O(d^{-1/2})$. In other words, it finds a ball whose volume is within $\exp(O_{\delta,\gamma}(\sqrt{d}))$ factor of the optimum, which is significantly stronger. Finally, the ball $\widehat{B}$ output by the algorithm also has coverage $(1-\gamma)\delta-O(\sqrt{d/n})$ by standard concentration arguments (see e.g., Devroye and Lugosi, 2001) since the VC-dimension of $d$-dimensional Euclidean balls is $d+1$.
All of our algorithmic results are distribution-free i.e., they do not make any assumptions on the distribution $\mathcal{D}$. While it may be impossible to learn the distribution or even estimate the density at a given point (these tasks often incur an exponential dependence on the dimension $d$ even for smooth distributions), our results show that one can obtain efficient algorithms that take polynomial (in $d$) time and find approximately optimal dense sets from $\mathcal{C}$ covering at least $\delta$ mass.

### v2

@@ Abstract
Our main result is an algorithm that finds a confidence set whose volume is $\exp(\tilde{O}(d^{1/2}))$ factor competitive with the optimal ball having the desired coverage. It is surprisingly simple and also extends to finding confidence sets competitive against unions of $k$ balls, and improved guarantees under additional assumptions. The algorithm is improper (it outputs an ellipsoid). Combined with our computational intractability result for proper learning balls within an $\exp(\tilde{O}(d^{1-o(1)}))$ approximation factor in volume, our results provide an interesting separation between proper and (improper) learning of confidence sets.
@@ 1 Introduction
where $\operatorname*{\mathrm{vol}}(S)$ denotes the volume or Lebesgue measure of the set $S$. The minimum of (1) is known as the generalized quantile function (Einmahl and Mason, 1992; Polonik, 1997), an important quantity in statistical inference that plays a central role in a number of problems including estimating density level sets (Garcia et al., 2003), distribution support estimation (Schölkopf et al., 2001) and conformal prediction (Gao et al., 2025). A set containing at least $\delta$, say $0.9$, probability mass generalizes the idea of confidence intervals in one-dimension to a high-dimensional setting.
Solving (1) leads to a number of interesting applications in statistical inference. On the one hand, finding a set that contains a large probability mass of the distribution can be useful as a high-dimensional analogue of support estimation (Schölkopf et al., 2001). On the other hand, as $\delta\to 0$, the maximizer of (1) can be used as a mode estimator  (Sager, 1978, 1979). Algorithms that learn confidence sets for arbitrary distributions can also be used as a tool for estimating the uncertainty in the predictions of black-box machine learning models, as in conformal prediction (Gammerman et al., 1998; Vovk et al., 2005). Finally, algorithms for this task have to be robust to contamination present in the data, as the remaining $(1-\delta)$ fraction of the data could be outliers. The formulation in (1) leads to a different notion of robustness that is non-parametric and may be of independent interest in robust estimation (Rousseeuw, 1984, 1985; Van Aelst and Rousseeuw, 2009). There is also a rich body of work on robust statistics, where the remaining $1-\delta$ fraction can be arbitrary outliers, but these works often make stronger model assumptions about the $\delta$ fraction of points that are “inliers”  (Huber, 1964; Diakonikolas and Kane, 2023). See Section 1.5 for applications in statistics like conformal prediction, goodness-of-fit, robust estimation and testing unimodality.
@@
Consider the setting where there exists a good confidence set belonging to a family $\mathcal{C}$ of finite VC-dimension. When independent samples from the distribution $\mathcal{D}$ are available, one can solve the empirical version of (1) with polynomial sample complexity and establish approximate volume optimality as long as the concept class $\mathcal{C}$ has finite VC-dimension (Polonik, 1999; Scott and Nowak, 2005). This is related to the (agnostic) PAC learning setting in classification where the confidence sets $S$ can be identified with classifiers belonging to the concept class $\mathcal{C}$. This leads to the following broad class of learning problems that we study:
@@
If there exists a small confidence set $C^{*}\in\mathcal{C}$ with (slightly more than) the required coverage $\delta$, then the algorithm will output a set $S$ with coverage $\delta$ and whose volume is competitive with $C^{*}$. The set $S$ is not required to belong to the family $\mathcal{C}$. This is analogous to the standard setting in PAC learning, where the learning algorithm may output a classifier that does not belong to the concept class $\mathcal{C}$ (see e.g., Shalev-Shwartz and Ben-David, 2014). If the algorithm outputs a confidence set $S$ that belongs to $\mathcal{C}$, we will call it a proper learner of confidence sets (analogous to proper PAC learning). When we want to be more explicit, the term improper learner may be used for an algorithm that does not necessarily output a set belonging to the concept class $\mathcal{C}$. We emphasize the worst-case nature of the distribution that the samples are drawn i.i.d from. This is analogous to the distribution-free setting of PAC learning, where we make no assumptions on the distribution. While there is a simple algorithm that is sample efficient (but computationally inefficient) when the VC-dimension of $\mathcal{C}$ is bounded (by a polynomial), the algorithmic problem becomes challenging in high dimensions.
@@
Consider the basic setting where $\mathcal{C}$ is the set of $\ell_{2}$ balls (this has VC-dimension at most $d$). This is a natural generalization of intervals in higher dimensions, and is motivated by applications (see e.g., Wang et al., 2023). It is NP-hard to find the smallest set in $\mathcal{C}$ that covers a $\delta$ fraction of points from an arbitrary set (Shenmaier, 2015). [^1] On the other hand, finding the minimum volume ball that encloses all the given $n$ points is polynomial time solvable. The famous work of Badoiu, Har-Peled and Indyk on coresets for minimum volume enclosing balls (Badoiu et al., 2002; Agarwal et al., 2004; Ding, 2020) gives a proper learning algorithm that runs in polynomial time and is $\Gamma=1+O(1/\log d)$-competitive i.e., a $\exp\big{(}O(d/\log d)\big{)}$ factor approximation in volume. To the best of our knowledge, these works using core-sets (Badoiu et al., 2002; Ding, 2020) remain the state-of-the-art guarantee for learning confidence sets competitive with Euclidean balls in $d$ dimensions. We focus on Euclidean balls and unions of Euclidean balls in the rest of the paper.
@@ 1.1 Our Results
Our main algorithmic result is an algorithm that is $\Gamma\approx 1+d^{-1/2}$ competitive against Euclidean balls.
@@
There is a polynomial time algorithm that for any target coverage $\delta\in(0,1)$ and coverage slack $\gamma\in(0,1)$, when given $n=\Omega(d^{2}/\gamma^{2})$ samples drawn i.i.d. from an arbitrary distribution $\mathcal{D}$, finds with high probability a set $S\subset\mathbb{R}^{d}$ that is $\Gamma=\exp\left(O_{\gamma,\delta}\big{(}d^{-1/2+o(1)}\big{)}\right)$ competitive, i.e.,
@@
$\operatorname*{\mathrm{vol}}(S)^{1/d}\leq\operatorname*{\mathrm{vol}}(B^{\star})^{1/d}\left(1+O\left(d^{-1/2+o(1)}(\gamma\delta)^{-1}\right)\right).$
@@
Please see Section 4 for more details. To compare against existing methods, consider the setting when $\delta,\gamma$ are constants. A straightforward solution that considers balls around the sample points that covers at least $\delta$ fraction of the data points achieves a factor of $\Gamma=2$ (due to the triangle inequality). [^2] In fact this strategy achieves a $\Gamma=2$ factor approximation (with $\textrm{vol}^{1/d}$) for any set $\mathcal{C}$ that consists of scalings and translations of a fixed convex shape. However, it is NP-hard to obtain a $(2-o(1))$ factor approximation even for the $\ell_{\infty}$ ball (Shenmaier, 2015). Algorithms based on the powerful technique of coresets for minimum volume enclosing balls due to  Badoiu et al. (2002) gives $\Gamma=1+\tilde{O}(1/\log d)$-competitive. In terms of the volume, the $\Gamma\approx 1+d^{-1/2}$ factor corresponds to volume approximation of
$\Gamma^{d}=\big{(}1+O_{\delta,\gamma}(d^{-1/2+o(1)})\big{)}^{d}=\exp\left(O_{\gamma,\delta}\big{(}d^{1/2+o(1)}\big{)}\right),$
which corresponds to a substantial improvement from the $\exp\big{(}O_{\delta,\gamma}(d/\log d)\big{)}$ volume approximation factor in existing work (Badoiu et al., 2002; Ding, 2020).
@@
Let $\delta\in(0,1),\gamma\in(0,1),k\in\mathbb{N}$ be any constants. There is a polynomial time algorithm that for target coverage $\delta\in(0,1)$ and coverage slack $\gamma\in(0,1)$ when given $n=\Omega(kd^{2}/\gamma^{2})$ samples drawn i.i.d. from an arbitrary distribution $\mathcal{D}$, finds with high probability a set $S\subset\mathbb{R}^{d}$ that is $\Gamma=\left(1+O_{\gamma,\delta}\big{(}d^{-1/2+o(1)}\big{)}\right)$ competitive; more precisely, it satisfies $\mathbf{P}_{y\sim\mathcal{D}}\left[y\in S\right]\geq\delta$, and
$\operatorname*{\mathrm{vol}}(S)^{1/d}\leq\mathrm{vol}(C^{\star}_{k})^{1/d}\left(1+O_{k,\delta}\left(d^{-1/2+o(1)}\right)\right)\cdot\left(\frac{O(\log(k/\gamma))}{\gamma}\right)^{1/d}$
@@ 1.2 Proper Learning: Separation Result and Algorithms
For any constant $\varepsilon>0$, assuming $P\neq NP$ there is no polynomial time learning algorithm that given samples from an arbitrary distribution over $\mathbb{R}^{d}$ that outputs balls as confidence sets that are $\Gamma\leq(1+\frac{1}{d^{\varepsilon}})$ competitive with balls.
See Theorem 6.1 for a more formal statement. Note that the above result is a worst-case hardness result, and reflects the arbitrary nature of the distribution. This is reminiscent of NP-hardness of proper learning in supervised learning (see e.g., Feldman et al., 2009). The above theorem follows from a simple reduction that amplifies the existing NP-hardness result due to (Shenmaier, 2015). In Theorem 6.3, we also show computational intractability (assuming the Small Set Expansion hypothesis of Raghavendra and Steurer (2010)) that shows a similar quantitative hardness, even when the coverage can be smaller by an arbitrary constant factor. Combined with our algorithmic result in Theorem 1.1, this shows a formal separation between proper and improper learning for a natural geometric learning problem.
@@
When we are required to output a confidence set from $\mathcal{C}$, the algorithm in Theorem 1.1 can also be used to output a ball with a worse competitive ratio (which is expected due to Theorem 1.3). When we make no assumptions about the distribution $\mathcal{D}$, our algorithm can output a ball that is $\Gamma=1+\widetilde{O}(1/\log d)$ competitive. See Theorem 3.1 in Section 3 for a formal statement of the theorem. This guarantee recovers the guarantee from the coreset-based algorithm which achieves $\Gamma=1+\widetilde{O}(1/\log d)$ (Badoiu et al., 2002). In the worst-case instances of the algorithm (and the hard instances in Theorem 1.3), the points sampled from $\mathcal{D}$ that lie inside the optimal ball $B^{\star}$ effectively lie on a lower-dimensional subspace.
@@
Let $Y\subseteq\mathbb{R}^{d}$ be a set of $n$ points, and $\delta,\gamma\in(0,1)$, such that there is an unknown ball $B^{\star}=B(c^{\star},R^{\star})$ with $Y^{\star}=B^{\star}\cap Y$ satisfying $|Y^{\star}|\geq\delta|Y|$, and $Y^{\star}$ is $\beta(R^{\star})^{2}/d$-isotropic for some $\beta\geq 0$. That is, if $\mu^{\star}$ is the mean of the points $Y^{\star}$, then
@@
$\operatorname*{\mathrm{vol}}(\widehat{B})^{1/d}\leq\operatorname*{\mathrm{vol}}(B^{\star})^{1/d}\sqrt{1+O\left(\frac{\beta}{\gamma\delta d}\right)},$
@@
In the above Theorem 1.4, $\Sigma_{Y^{\star}}$ is a PSD matrix of dimension $d\times d$ with $\mathrm{tr}(\Sigma_{Y^{\star}})\leq(R^{\star})^{2}$, since $Y^{\star}\subseteq B^{\star}$. The isotropicity condition requires the maximum eigenvalue $\lambda_{\max}(\Sigma_{Y^{\star}})\leq\beta(R^{\star})^{2}/d$ for some $\beta>0$. Thus every $Y^{\star}$ satisfies the isotropic condition with $\beta\leq d$. When the points of $Y^{\star}$ are very spread out, this corresponds to a setting when $\beta\leq 1$ (all the eigenvalues are equal). In this case, the theorem guarantees to find a ball with $\Gamma=1+O(d^{-1})$. In other words, it finds a ball whose volume is within $O_{\delta,\gamma}(1)$ factor of the optimum, which is significantly stronger. Finally, the ball $\widehat{B}$ output by the algorithm also has coverage $(1-\gamma)\delta-O(\sqrt{d/n})$ by standard concentration arguments (see e.g., Devroye and Lugosi, 2001) since the VC-dimension of $d$-dimensional Euclidean balls is $d+1$.
All of our algorithmic results are distribution-free i.e., they do not make any assumptions on the distribution $\mathcal{D}$. While it may be impossible to learn the distribution or even estimate the density at a given point (these tasks often incur an exponential dependence on the dimension $d$ even for smooth distributions), our results show that one can obtain polynomial time algorithms that find approximately optimal dense sets from $\mathcal{C}$ covering at least $\delta$ mass.